### v1

# SPARC: Prediction-Based Safe Control for Coupled Controllable and Uncontrollable Agents with Conformal Predictions

Abstract
We investigate the problem of safe control synthesis for systems operating in environments with uncontrollable agents whose dynamics are unknown but coupled with those of the controlled system. This scenario naturally arises in various applications, such as autonomous driving and human-robot collaboration, where the behavior of uncontrollable agents, like pedestrians, cannot be directly controlled but is influenced by the actions of the autonomous vehicle or robot. In this paper, we present SPARC (Safe Prediction-Based Robust Controller for Coupled Agents), a novel framework designed to ensure safe control in the presence of coupled uncontrollable agents. SPARC leverages conformal prediction to quantify uncertainty in data-driven prediction of agent behavior. Particularly, we introduce a joint distribution-based approach to account for the coupled dynamics of the controlled system and uncontrollable agents. By integrating the control barrier function (CBF) technique, SPARC provides provable safety guarantees at a high confidence level. We illustrate our framework with a case study involving an autonomous driving scenario with walking pedestrians.
I INTRODUCTION
Autonomous systems are increasingly being deployed in our society with a wide range of applications, from self-driving cars to industrial robots. Ensuring safety is one of the most critical tasks in these systems, as they must operate in dynamic environments while avoiding collisions and other unsafe behaviors. For example, in autonomous driving, the system must continuously assess its surroundings to ensure the safety of pedestrians, other vehicles, and passengers.
To achieve safe control in real-world scenarios, one must take the presence of uncontrollable agents in the environment into account. These agents, such as pedestrians or other vehicles, move and make decisions independently of the autonomous system. For instance, in an urban driving scenario, a self-driving car must account for the unpredictable behavior or intention of pedestrians, who may change speed or direction in response to the car’s actions. Managing this interaction is crucial to ensuring the overall safety of the system.
There are two major challenges in achieving safe control in the presence of uncontrollable agents. First, the dynamics of the controlled system and the uncontrollable agents are generally coupled. In other words, the behavior or intention of the uncontrollable agent depends on the state of the controlled system. For example, in autonomous driving, a pedestrian’s movements may change in response to the position and speed of a nearby vehicle. Second, the intentions of such uncontrollable agent are often unknown a priori or very difficult to model precisely. As a result, traditional model-based approaches, such as robust control, cannot be reliably used to ensure safety. The inability to predict the agent’s exact behavior makes it challenging to design a control strategy that accounts for all possible interactions with the uncontrollable agent.
Data-driven approaches provide a powerful tool for handling the unknown intentions of uncontrollable agents. By leveraging large datasets collected from real-world interactions, one can develop highly accurate predictors to forecast the trajectories of uncontrollable agents for control purposes. For instance, as shown in Figure 1, in autonomous driving within pedestrian-rich environments, there are extensive datasets that capture how pedestrians react to vehicles under various conditions. However, a common drawback of data-driven approaches is the lack of formal guarantees. While the prediction of a pedestrian’s intention may be correct 95% of the time, the remaining 5% of failures can lead to unacceptable risks. In safety-critical systems, such errors are intolerable, and one has to design a robust system by taking the uncertainty in data-driven approaches into account.
(a) Intersections Drone Dataset (inD) [1]
(b) Stanford Drone Dataset (SDD) [2]
Figure 1: Instances of coupled dynamics in real-world scenarios can be observed across various datasets. For example, the interactions between traffic participants often exhibit mutual influence. In Fig. 1(a), vehicles may decelerate as they encounter bicycles at crosswalks. At the same time bicycles may alter their trajectories to avoid collisions with vehicles, resulting in curved paths.
In this paper, we propose a new framework, called SPARC (Safe Prediction-Based Robust Controller for Coupled Agents), for ensuring safe control in the presence of uncontrollable agents. Specifically, we assume the dynamics of the uncontrollable agent are unknown and coupled with the dynamics of the controlled system. To address this challenge, we leverage the theory of conformal prediction to quantify uncertainty in data-driven predictions of the uncontrollable agent’s intentions. Unlike previous approaches, which assume that the dynamics of the uncontrollable agent are independent of the controlled system, we develop a novel method to sample data and construct conformal regions based on the joint distribution over an augmented state space. This approach allows us to capture the inter-dependencies between the two agents and improve prediction accuracy. Furthermore, by incorporating the control barrier function (CBF) technique, we can provide provable safety guarantees with a confidence bound, ensuring robust system performance. Our framework is particularly well-suited for applications such as autonomous driving in urban environments or human-robot collaborations, where the uncontrollable environments have their own underlying decision-laws that depend on the behavior of the controlled systems.
Figure 2: SPARC (Our Framework)
II Related Works
Safety-Critical Control using CBFs. Control Barrier Functions (CBFs) are a widely used approach for ensuring safety in control systems [3, 4, 5, 6, 7]. First introduced by Ames et al. [8], the key idea behind CBFs is to encode safety constraints into a function that guarantees the system remains within a safe set during operation. However, a significant challenge with CBFs is their reliance on an accurate dynamic model of the system. When the system includes an uncontrollable component with unknown dynamics, applying traditional CBFs can be problematic. In such cases, robust CBFs [9, 10] can be used to account for the worst-case scenario, which, however, often results in overly conservative control strategies.
Data-Driven Safety Control Synthesis. When the system dynamic is unknown a priori, many recent works have studied how to synthesize safe controllers using data; see, e.g., [11, 12, 13, 14, 15, 16, 17]. For instance, in [15], the authors convert the safe control synthesis to a convex program by using barrier certificates. The scenario approach is used to provide a probabilistic safety guarantee based on sampled data. However, most of these works cannot handle the presence of uncontrollable agents whose dynamics are coupled with the control systems.
Conformal Prediction in Control Synthesis. Conformal prediction is an emerging technique in machine learning that provides a way to quantify the uncertainty of model predictions [18]. It works by generating prediction sets or confidence regions that, with a specified level of certainty, are guaranteed to contain the true outcome. This technique is model-free, meaning it can be applied to any underlying predictive model, and it offers formal statistical guarantees on the accuracy of its predictions. Recently, conformal prediction has been extensively applied to control synthesis problems; see, e.g., [19, 20, 21, 22, 23, 24]. Our work is closely related to [22] and [19], which address the problem of safe planning in unknown dynamic environments. In these works, the authors also employ conformal prediction to estimate the confidence region of the uncontrollable agent. However, a key distinction is that these approaches assume the behavior of uncontrollable agents is independent of the controlled system, an assumption that may not hold in many real-world scenarios. In contrast, our framework considers a coupled setting, addressing the issue of distribution shift in conformal prediction, where the behavior of uncontrollable agents is influenced by the state of the controlled system.
III Problem Formulation
Notations: We denote by $\mathbb{R},\mathbb{N}$, and $\mathbb{R}^{n}$ the set of real numbers, natural numbers, and real vectors, respectively. Let $\beta:\mathbb{R}\rightarrow\mathbb{R}$ denote an extended class $\mathcal{K}_{\infty}$ function, i.e., a strictly increasing function with $\beta(0)=0$. For a vector $v\in\mathbb{R}^{n}$, let $\|v\|$ denote its Euclidean norm. For a bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, let $\partial\mathcal{S}$ denote its boundary.
III-A System Models
Controlled System: The system under control is modeled as a nonlinear control-affine system of form
$\dot{x}(t)=f(x(t))+g(x(t))u(t)=:F(x(t),u(t)),$ (1)
where $x(t)\in\mathbb{R}^{n}$ represents the state of the system at time $t$, and $u(t)\in\mathcal{U}$ denotes the control input, with $\mathcal{U}\subseteq\mathbb{R}^{m}$ being the set of admissible control inputs. The functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ represent the internal system dynamics and the input influence, respectively, and both are assumed to be locally Lipschitz continuous. For convenience, we denote $\mathbf{X}(t)=[x(t),\dot{x}(t)]\in\mathbb{R}^{n_{X}}$.
Uncontrollable Agent: We consider the scenario where the controlled system in Eq. (1) operates among some uncontrollable agents. For the sake of simplicity, we only consider a single uncontrollable agent, and our approach can be extended to $N$ uncontrollable agents by considering their states as a vector. We denote the state of the agent at time instant $t$ by $Y(t)$, where $Y(t)\in\mathbb{R}^{n_{Y}}$. Note that $n_{Y}$ does not necessarily equal $n$, since the dynamics of $X$ and $Y$ may differ.
More specifically, we assume that the uncontrollable agent has its own dynamic and decision law (intention), which can be either deterministic or stochastic. However, we do not assume to have any a priori knowledge of the agent dynamics or intentions. Note that, in general, the behavior of the uncontrollable agent depends on the state of the controlled agent. For example, in autonomous driving scenarios with moving pedestrians (uncontrollable agents), the behavior of the pedestrians is significantly influenced by the position and velocity of the car. Specifically, the closer the car is, the faster pedestrians tend to move away, or the faster the car approaches, the more cautious pedestrians become in their walking strategies.
The objective is to synthesize a control input $u(t)$ for each time instant such that the controlled system can ensure safety with respect to the uncontrollable agent, with a certain level of confidence. Formally, safety here can be framed as a safe set $\mathcal{S}\subseteq\mathbb{R}^{n}\times\mathbb{R}^{n_{Y}}$. Then the problem formulation is as follows.
Problem 1
Consider a controlled system described by Eq. (1) and an uncontrollable agent whose behavior depends on the behavior of the controlled system. Suppose that the controller can access the states of both the controlled system and uncontrollable agent precisely without measurement error, but the physical dynamics and intentions of the uncontrollable agent is unknown a priori. Design a controller that issues control input $u(t)$ for each time instant such that the closed-loop behavior is safe w.r.t. safety set $\mathcal{S}$, with a certain level of confidence.
Since the behavior of the uncontrollable agent is complete unknown, hereafter, we will use sample data to provide confidence level guarantee for safety of form: “the probability that the system becomes unsafe within a certain time horizon is smaller than a threshold bound $\alpha$”.
IV Safe Prediction-Based Robust Controller for Coupled Agents (SPARC)
IV-A Overall Control Framework
Before we formally present the implementation details of the proposed controller, we provide a high level overview of the entire SPARC framework.
  • At each discrete time step, referred to as a prediction update point, we use a neural network to predict the intention of the uncontrollable agent over the next control period $\Delta t$.
  • We apply the technique of conformal predictions to enlarge the predicted results into a region that contains the agents’ reachable states with high confidence.
  • Then within each control period $\Delta t$, we apply the technique of control barrier functions to generate control inputs continuously to avoid reaching the predicted region of the uncontrollable agent.
  • The above process is repeated for prediction update point indefinitely.
To implement the generic idea of the above framework, the main challenge lies in the coupling between the controlled system and the uncontrollable agent. Specifically, the behavior of the uncontrollable agent depends on the behavior of the controlled system. Therefore, one cannot simply design a predictor and apply conformal prediction solely based on the trajectory data of the uncontrollable agent, as the agent will behave differently when the controlled system is in a different state. This issue is the well-known distribution shift problem in conformal prediction. To address this challenge, our approach is to design a predictor and perform conformal prediction over an augmented state space that accounts for the joint behavior of both the controlled system and the uncontrollable agents.
IV-B Data Collection and Predictor Design
To collect data and to predict the behavior of the uncontrollable agent, we sample the trajectories of both the controlled system and the uncontrollable agent with time interval $\Delta t$. We denote by $[X_{0},X_{1},\ldots,X_{k}]$ and $[Y_{0},Y_{1},\ldots,Y_{k}]$ the sampled trajectories. To capture the coupling between the controlled system and the uncontrollable agents, at each sampling instant, the movement of the uncontrollable agent, i.e., $\Delta Y_{i+1}:=Y_{i+1}-Y_{i}$, follows a distribution parameterized by the state of the controlled system, i.e.,
$\Delta Y_{i+1}\sim\mathcal{D}(X_{i}),.$ (2)
Note that, distribution $\mathcal{D}(X_{i})$ is memoryless and time-invariant meaning that the behavior of the uncontrollable agent only depends on the current state of the system rather than the entire history. This assumption is valid in many applications such as the behavior of moving pedestrians. Furthermore, this distribution is unknown a priori and we can only sample data according to the distribution.
Data Collections.  We assume that we have access to $\bar{K}$ independent trajectories of observations $X^{(j)}=(X_{1}^{(j)},X_{2}^{(j)},\ldots)$, $Y^{(j)}=(Y_{1}^{(j)},Y_{2}^{(j)},\ldots)$. Let $\Delta Y_{i+1}=Y_{i+1}-Y_{i}$, $\Delta Y^{(j)}=(\Delta Y_{1}^{(j)},\Delta Y_{2}^{(j)},\ldots),\forall j=1,2,\ldots,\bar{K}$ sampled from the distribution of $X^{(j)}$. Let $Z^{(j)}=(Z_{1}^{(j)},Z_{2}^{(j)},\ldots)$ where $Z_{i}^{(j)}=(X_{i}^{(j)},\Delta Y_{i}^{(j)})$. We partition these trajectories into calibration and training sets, denoted as $D_{\text{cal}}:=\{Z^{(1)},\dots,Z^{(Card(D_{cal}))}\}$ and $D_{\text{train}}:=\{Z^{(Card(D_{cal})+1)},\dots,Z^{(\bar{K})}\}$, respectively.
Trajectory Predictor.  Our objective is to use the training set $D_{\text{train}}$ to design a predictor to forecast the future state of the uncontrollable agent. Note that, since $\Delta Y_{i+1}+Y_{i}=Y_{i+1}$, it suffices to predict $\Delta Y_{i+1}$ as we assume that all state information can be observed perfectly. In principle, although the $\Delta Y_{i+1}$ only depends on the current states, we still use the entire past samples to perform the prediction to improve the prediction accuracy. Therefore, the predictor is of form
$\Delta\hat{Y}_{t+1}=\Omega(Z_{(0,t]}|\theta).$ (3)
Note that $\Omega$ could be implemented by any trajectory predictor such as recurrent neural networks (RNNs), long short term memory (LSTM) networks or transformers , and $\theta$ is its parameters.
IV-C Conformal Prediction for Reachable Region
The predictions $\Delta\hat{Y}_{t}$, which may be generated by models with learning capabilities, are not guaranteed to be accurate. To address this, we employ conformal prediction to construct prediction regions for $\Delta\hat{Y}_{t}$ that hold with high probability.
Lemma 1
([25]). Let $V,V^{(1)},\ldots,V^{(k)}$ be $k+1$ independent and identically distributed real-valued random variables. Without loss of generality, let $V^{(1)},\ldots,V^{(k)}$ be sorted in non-decreasing order and define $V^{(k+1)}:=\infty$. For $\alpha\in(0,1)$, it holds that $\operatorname{Prob}(V\leq\bar{V})\geq 1-\alpha$ where
$\bar{V}:=V^{(r)}\text{ with }r:=\lceil(k+1)(1-\alpha)\rceil$
[…]
In the context of our approach, we aim to use conformal prediction to quantify the uncertainty in the predicted state of the uncontrollable agent.
Definition 1
For a trajectory $Z^{(j)}=(Z_{1}^{(j)},Z_{2}^{(j)},\ldots)\in D_{cal}$ and $Z^{(j)}_{i}=(X^{(j)}_{i},\Delta Y^{(j)}_{i})$, the prediction error is then defined as
$R^{(j)}:=\sup_{Z^{(j)}_{i}\in Z^{(j)}}||\Delta Y_{i}^{(j)}-\Delta\hat{Y}_{i}^{(j)}||.$ (4)
$\Delta Y_{i}^{(j)}$ is regarded as the ground truth of $\Delta\hat{Y}_{i}^{(j)}$.
The following result states how to use the calibration dataset to generate a confidence regions that account for potential variations in the agent’s behavior.
Proposition 1
Given a calibration dataset $D_{cal}\subseteq\mathbb{R}^{(n_{X},n_{Y}})$, and a failure probability tolerance $\alpha\in[0,1]$, by Lemma 1, $\forall X\in\mathbb{R}^{n_{X}}$, $t\in\mathbb{R}_{\geq 0}$, $\operatorname{Prob}\left(R^{(0)}\leq\bar{R}(\alpha)\right)\geq 1-\alpha$ holds where
$\bar{R}(\alpha)=\text{ Quantile }_{1-\alpha}\left(R^{(1)},\ldots,R^{(Card(D_{cal}))},\infty\right)$ (5)
Note that here $\operatorname{Card}(D_{\text{cal}})$ is the dataset size as in Assumption 1. Similarly, we obtain $\bar{R}=R^{(r)}$, where $r=\lceil\operatorname{Card}(D_{\text{cal}})\cdot(1-\alpha)\rceil$. As before, for meaningful prediction regions, we require $\operatorname{Card}(D_{\text{cal}})\geq\lceil(\operatorname{Card}(D_{\text{cal}})+1)(1-\alpha)\rceil$, otherwise $\bar{R}(\alpha)=\infty$.
Proof:
In Lemma 1, given a set of independent and identically distributed random variables $V^{(1)},V^{(2)},\ldots,V^{(k)}$, we can derive a bound in a sense of probability, by sorting these values and determining the corresponding quantile. In Proposition 1, we treat the nonconformity scores $R^{(1)},R^{(2)},\ldots,R^{(\text{Card}(D_{\text{cal}}))}$ as analogous to the random variables $V^{(1)},V^{(2)},\ldots,V^{(k)}$ in Lemma 1. Therefore, $R$ plays the role of $V$ in the lemma, allowing us to directly apply its result. ∎
The above result says that the state estimation error $R$ can essentially be bounded, with probability $1-\alpha$, by the worst case of conformal prediction region $\bar{R}(\alpha)$. Since the score $R^{(0)}$ represents the model’s estimation error sampled from any trajectory that might occur during the actual operation of the system, we can hence conclude that
$\operatorname{Prob}[||\Delta Y_{t+1}-\Delta\hat{Y}_{t+1}||\leq\bar{R}(\alpha)]\geq 1-\alpha$ (6)
From known measurement $X_{t}$ and $Y_{t}$, and prediction $\Delta\hat{Y}_{t+1}$, denote safe region in a probabilistic sense for $\Delta\hat{Y}_{t}$ and $\bar{R}(\alpha)$ as $\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)=[(Y_{t}+\Delta\hat{Y}_{t+1})-\bar{R}(\alpha),(Y_{t}+\Delta\hat{Y}_{t+1})+\bar{R}(\alpha)]\subseteq\mathbb{R}^{n_{Y}}$ satisfies that
$\operatorname{Prob}[Y_{t+1}\in\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)]\geq 1-\alpha$ (7)
Note that $\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)$ is the estimated safe region of the uncontrollable event within the next sample period.
IV-D Safe Control using Control Barrier Functions
After bounding the state estimation error in Proposition 1, we calculated the conformal prediction region bound $\bar{R}$ for Eq. (6) to design a uncertainty-aware controller via the technique of control barrier functions.
Definition 2
([8]) Let $\mathcal{C}\subseteq\mathbb{R}^{n}$ be the zero-superlevel set of a continuously differentiable function $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$. The function $h$ is a control barrier function (CBF) for the system in Eq. (1) if there exists an extended class $\mathcal{K}_{\infty}$ function $\beta$ such that
$\sup_{u\in\mathcal{U}}[L_{f}h(x)+L_{g}h(x)]\geq-\beta(h(x))$ (8)
where $L_{f}h(x)$ and $L_{g}h(x)$ represent the Lie derivatives.
Therefore, Problem 1 can be formulated as a safe control problem within each prediction update period by considering the estimated region of the uncontrollable agent as an obstacle simply by define
$\mathcal{C}=\{x\in\mathbb{R}^{n}\mid\forall Y\in\mathbb{R}^{n_{Y}}:(x,Y)\in\mathcal{S}\}$ (9)
as the safe set for the controlled system such that the overall system remain in safe set $\mathcal{S}$. Therefore, it suffices to find a CBF $\hat{h}(x)$ such that
$\hat{C}=\{x\in\mathbb{R}^{n}\mid\hat{h}(x)\geq 0\}\subseteq\mathcal{C}.$ (10)
In the literature, there are many existing approach for finding such CBF $\hat{h}(x)$, e.g., by sum-of-square programming [26] or learning-based approach [27, 28].
Now, given the synthesized CBF $\hat{h}(x)$, the set of admissible control inputs is defined by
$K_{CBF}:=\{u\in\mathcal{U}\mid L_{f}\hat{h}(x)+L_{g}\hat{h}(x)u\geq-\beta(\hat{h}(x))\}$ (11)
The following result shows that, by following the control strategy in Eq. (11), one can ensure safety within each period $\Delta$ with confidence $1-\alpha$.
Proposition 2
Consider the control law $u(t)$ from $K_{CBF}$ in Eq. (11),the system composed by controllable and uncontrollable variables Eq. (1), Eq. (2) satisfies Eq. (8) where $\mathcal{L}_{L_{f}h},\mathcal{L}_{\beta\circ h}$, and $\mathcal{L}_{L_{g}h}$ are the Lipschitz constants of the functions $L_{f}h,\beta\circ h$ and $L_{g}h$, respectively, we have
$\operatorname{Prob}[(X_{t+1},Y_{t+1})\in\mathcal{S}]\geq 1-\alpha$ (12)
Proof:
When $Y_{t+1}\in\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)$ (with probability higher than $1-\alpha$), the“obstacle”(i.e., the safe region which includes the actual position of the uncontrollable agent inside) is accurately estimated. Consequently, the control law $u(t)$ derived from $K_{CBF}$ for $\hat{\mathcal{C}}$ is effective in ensuring the safety of the entire system, meaning that $(X_{t+1},Y_{t+1})$ remains within the safe set $\mathcal{S}$, as defined in Problem 1. ∎
V Simulation Results
In this section, we illustrate and evaluate the proposed approach through simulations of a case study involving autonomous driving with uncontrollable pedestrians. This case study demonstrates the effectiveness of our framework in managing the interactions between the autonomous vehicle and pedestrians, ensuring safety despite the uncertainty in pedestrian behavior.
Figure 3: The pedestrian’s walking logic is modeled as an unknown stochastic distribution affected by the state of the vehicle, i.e.,$\Delta Y\sim\mathcal{D}(X)$. The pedestrian starts at a random position on one side of the crosswalk while the vehicle begins at the midpoint of the starting line. Red circles indicate collision volume.
Figure 4: Nonconformity scores histogram for pedestrian position displacement $(\Delta x,\Delta y)$ on $\mathcal{D}_{cal}$
V-A Scenario Description and System Dynamic
Scenario Setup: As illustrated in Fig. 3, the scenario is set on a straight road segment with a total length of 84 m. A crosswalk is located within this segment, and the road has a width of 25 m. The starting line for the vehicle is positioned at one end of the road, while the ending line is at the opposite end. A pedestrian is present near the crosswalk, and a vehicle is approaching from the starting line. The control period is set to be $\Delta t$ = 0.1s.
Vehicle Dynamics: The vehicle’s starting speed is sampled from a uniform distribution within the range $[0,15]$ m/s, which mimics typical vehicle speeds observed in urban environments (up to the speed limit of 15 m/s). The vehicle’s position is updated based on its velocity.
Pedestrian Behavior: The pedestrian’s movement is modeled as a unknown distribution of the vehicle’s state-space vector and Gaussian noise i.e.
$\Delta{Y}=\phi(X)+\mathcal{N}({0,\sigma})\sim\mathcal{D}(X)$ (13)
[…]
Interaction and Collision Avoidance: The pedestrian’s trajectory is evaluated in real-time as the vehicle approaches. We track the pedestrian’s crossing behavior when the pedestrian is on the crosswalk, where the vehicle’s speed directly impacts the pedestrian’s decision to speed up, slow down, or stop. The simulation continuously monitors the distance between the vehicle and the pedestrian, and the safe controller(SPARC) is applied to ensure safe control synthesis in avoiding collisions.
Trajectory Predictor: We trained a feedforward neural network to predict the future trajectory of the pedestrian. The input is the history state measurement of the vehicle’s and pedestrian’s state. The training dataset $D_{\text{train }}$ contains $1\times 10^{6}$ data points, and the calibration dataset $D_{\text{cal }}$ for conformal prediction contains $1\times 10^{5}$ data points.
Barrier Functions: For vehicle-pedestrian collision avoidance, we may define the CBF as follows:
$h(x)=\min_{Y\in\partial\hat{\mathcal{O}}_{Y}}||x-Y||-d_{\text{safe}}$ (14)
where:
  • $x\in\mathbb{R}^{n}$ is the position of the vehicle
  • $\partial\hat{\mathcal{O}}_{Y}\subseteq\mathbb{R}^{n_{Y}}$ is the predicted safe region boundary of the uncontrollable pedestrian
  • $d_{\text{safe}}$ is the minimum safe distance to be maintained
[…]
This CBF in a sense of probability ensures a safe distance is maintained between the vehicle and the pedestrian. The system is in a safe state, avoiding collision, when $h(x)\geq 0$. The controller design should satisfy the inequality in Eq. (8) to guarantee system safety. Set $d_{\text{safe}}$ = 1.0 m, the collision volume is shown as the red circle in Fig. 3.
To demonstrate the effectiveness of our method, we compare the following two cases in simulations:
  1. 1.
    Uncontrollable Prediction Robust CBF: Let $\alpha$ be the failure tolerance, randomly pick $\alpha=0.85,0.75,0.5$, the results are shown in Table I.
  2. 2.
    Vanilla CBF: We set $\alpha=0\implies\bar{R}(\alpha)=[0,0,\ldots,0]$, which essentially reduces to the vanilla CBF [8]. The barrier function is directly from neural network output rather than safe region, so this CBF cannot provide any safety guarantee.
V-B Result Analysis
Figure 5: The result of the conformal prediction in the case study
Trajectory Predictors and Conformal Prediction: Take failure probability tolerance $\alpha=0.85$ as an example, Fig. 4 illustrates a calculation of the nonconformity scores used in our approach. The scores $R_{x}^{0.85}=0.75,R_{y}^{0.85}=2.86$ represent the deviation between the predicted pedestrian position change $\Delta\hat{Y}=(\Delta x,\Delta y)$ and the ground truth, providing a measure of prediction uncertainty , which is used to generate safe control decisions.
Fig. 5 shows the network successfully learns the underlying random distribution governing pedestrian behavior in relation to the vehicle’s state. In our simulation framework, for the sake of simplicity, the upper limit of all nonconformity scores is employed as a proxy for the overall safety margin, and it is also feasible to utilize a grid-based distribution of nonconformity scores to devise a more aggressive control strategy.
Safe Control Barrier Function Controller: The vehicle driven by our SPARC is expected to avoid collisions with pedestrians. We set the failure probability tolerance $\alpha=0.15,0.25,0.5$, and desire $\text{Prob}[(X,Y)\in\mathcal{S}]\geq 1-\alpha$, which means $\text{Prob}[\text{vehicle and pedestrian collide}]\leq\alpha$. We performed $1\times 10^{5}$ experiments, each time step is set to 100, the collision probability are collected in Table I. In 11.20% of the trials(1120 experiments), the car encounter a pedestrian, and we care about whether our controller can successfully prevent collision in these experiment, so we divide absolute collision frequency by 11.20% to get relative collision frequency. The collision rate of SPARC are $14.46\%(<15\%),17.50\%(<25\%),37.59\%(<50\%)$, which are all under their respective failure probability tolerance and significantly lower than vanilla CBF case $(96.43\%)$, so our Proposition 2. holds in practice.
TABLE I: Collision Probability
Controller Collision(Absolute) Collision(Relative)
SPARC ($\alpha$=0.15) 1.62% 14.46% < 15%
SPARC ($\alpha$=0.25) 1.96% 17.50% < 25%
SPARC ($\alpha$=0.50) 4.21% 37.59% < 50%
Vanilla CBF 10.80% 96.43%
Random 11.20% 100.00%
VI Conclusion
This work presents a new framework that integrates trajectory prediction with conformal prediction to enhance safety in multi-agent systems involving coupled controllable and uncontrollable agents. By incorporating probabilistic guarantees into the control barrier function framework, we ensure safe interactions between the controlled system and uncontrollable agents with high confidence. Compared to previous approaches, our method accounts for the coupling between the controlled system and the uncontrollable agent by sampling and evaluating over an augmented state space, making it more applicable to real-world scenarios such as autonomous driving and human-robot interaction. However, one limitation of our approach is the assumption that precise state information for both the controlled system and uncontrollable agent is fully accessible. In many practical applications, such state information must be obtained through perception devices like LiDAR, which can introduce estimation errors. In the future, we plan to extend our framework to perception-based settings, addressing the additional challenge of state estimation uncertainty.

### v3

@@
 Abstract
-We investigate the problem of safe control synthesis for systems operating in environments with uncontrollable agents whose dynamics are unknown but coupled with those of the controlled system.
-This scenario naturally arises in various applications, such as autonomous driving and human-robot collaboration, where the behavior of uncontrollable agents, like pedestrians, cannot be directly controlled but is influenced by the actions of the autonomous vehicle or robot.
-In this paper, we present SPARC (Safe Prediction-Based Robust Controller for Coupled Agents), a novel framework designed to ensure safe control in the presence of coupled uncontrollable agents.
-SPARC leverages conformal prediction to quantify uncertainty in data-driven prediction of agent behavior.
+We investigate the problem of safe control synthesis for autonomous systems operating in environments with uncontrollable agents whose dynamics are unknown but coupled with those of the controlled system.
+This scenario naturally arises in various applications, such as autonomous driving and human-robot collaboration, where the behavior of uncontrollable agents, such as pedestrians or human operators, cannot be directly controlled but is influenced by the actions or the states of the autonomous vehicle or robot.
+In this paper, we present SPARC (Safe Prediction-Based Robust Controller for Coupled Agents), a new framework designed to ensure safe control in the presence of uncontrollable agents that are coupled with the controlled system.
 Particularly, we introduce a joint distribution-based approach to account for the coupled dynamics of the controlled system and uncontrollable agents.
-By integrating the control barrier function (CBF) technique, SPARC provides provable safety guarantees at a high confidence level.
-We illustrate our framework with a case study involving an autonomous driving scenario with walking pedestrians.
+Then, SPARC leverages conformal prediction to quantify the uncertainty in data-driven prediction of the agent behavior.
+By integrating the control barrier function (CBF), SPARC provides provable safety guarantees at a high confidence level.
+We illustrate our framework with a case study of an autonomous driving scenario with walking pedestrians.
 I INTRODUCTION
-Autonomous systems are increasingly being deployed in our society with a wide range of applications, from self-driving cars to industrial robots.
-Ensuring safety is one of the most critical tasks in these systems, as they must operate in dynamic environments while avoiding collisions and other unsafe behaviors.
-For example, in autonomous driving, the system must continuously assess its surroundings to ensure the safety of pedestrians, other vehicles, and passengers.
-To achieve safe control in real-world scenarios, one must take the presence of uncontrollable agents in the environment into account.
-These agents, such as pedestrians or other vehicles, move and make decisions independently of the autonomous system.
-For instance, in an urban driving scenario, a self-driving car must account for the unpredictable behavior or intention of pedestrians, who may change speed or direction in response to the car’s actions.
-Managing this interaction is crucial to ensuring the overall safety of the system.
-There are two major challenges in achieving safe control in the presence of uncontrollable agents.
-First, the dynamics of the controlled system and the uncontrollable agents are generally coupled.
-In other words, the behavior or intention of the uncontrollable agent depends on the state of the controlled system.
+Safety is one of the major concerns in autonomous systems such as self-driving cars or industrial robots.
+These systems must operate in dynamic environments while avoiding unsafe behaviors.
+For instance, in autonomous driving, the ego vehicle must continuously monitor its surroundings to ensure the safety of itself, pedestrians, and other vehicles.
+To achieve safe control in real-world scenarios, it is essential to account for the presence of uncontrollable agents in the environment.
+There are two major challenges in achieving safe control in such scenarios.
+First, the dynamics of the controlled system and the uncontrollable agents are coupled in general.
+In other words, the behavior of the uncontrollable agent depends on the state of the controlled system.
 For example, in autonomous driving, a pedestrian’s movements may change in response to the position and speed of a nearby vehicle.
-Second, the intentions of such uncontrollable agent are often unknown a priori or very difficult to model precisely.
-As a result, traditional model-based approaches, such as robust control, cannot be reliably used to ensure safety.
-The inability to predict the agent’s exact behavior makes it challenging to design a control strategy that accounts for all possible interactions with the uncontrollable agent.
-Data-driven approaches provide a powerful tool for handling the unknown intentions of uncontrollable agents.
+Second, the intentions of such agents are often unknown a priori and difficult to model precisely.
+As a result, data-driven approaches have emerged as a powerful tool for handling the intentions of uncontrollable agents.
 By leveraging large datasets collected from real-world interactions, one can develop highly accurate predictors to forecast the trajectories of uncontrollable agents for control purposes.
 For instance, as shown in Figure 1, in autonomous driving within pedestrian-rich environments, there are extensive datasets that capture how pedestrians react to vehicles under various conditions.
-However, a common drawback of data-driven approaches is the lack of formal guarantees.
-While the prediction of a pedestrian’s intention may be correct 95% of the time, the remaining 5% of failures can lead to unacceptable risks.
-In safety-critical systems, such errors are intolerable, and one has to design a robust system by taking the uncertainty in data-driven approaches into account.
-(a) Intersections Drone Dataset (inD) [1]
-(b) Stanford Drone Dataset (SDD) [2]
+However, these predictors fail to measure uncertainty in a formal and systematic manner, rendering the prediction-based control susceptible to unsafe outcomes.
+(a) InD Dataset[1]
+(b) Stanford Drone Dataset[2]
 Figure 1: Instances of coupled dynamics in real-world scenarios can be observed across various datasets.
 For example, the interactions between traffic participants often exhibit mutual influence.
-In Fig.
-1(a), vehicles may decelerate as they encounter bicycles at crosswalks.
+In Fig.1(a), vehicles may decelerate as they encounter bicycles at crosswalks.
 At the same time bicycles may alter their trajectories to avoid collisions with vehicles, resulting in curved paths.
@@
 This approach allows us to capture the inter-dependencies between the two agents and improve prediction accuracy.
-Furthermore, by incorporating the control barrier function (CBF) technique, we can provide provable safety guarantees with a confidence bound, ensuring robust system performance.
-Our framework is particularly well-suited for applications such as autonomous driving in urban environments or human-robot collaborations, where the uncontrollable environments have their own underlying decision-laws that depend on the behavior of the controlled systems.
+Furthermore, by incorporating the control barrier function (CBF) techniques, we can provide provable safety guarantees with a confidence bound, ensuring robust system performance.
+Our framework is particularly well-suited for applications such as autonomous driving in urban environments or human-robot collaboration, where uncontrollable agents operate based on their own underlying hidden decision-making processes that are influenced by the behavior of the controlled systems.
 Figure 2: SPARC (Our Framework)
@@
 [8], the key idea behind CBFs is to encode safety constraints into a function that guarantees the system remains within a safe set during operation.
-However, a significant challenge with CBFs is their reliance on an accurate dynamic model of the system.
+However, a significant challenge of using CBFs is their reliance on an accurate dynamic model of the system.
 When the system includes an uncontrollable component with unknown dynamics, applying traditional CBFs can be problematic.
@@
 Conformal Prediction in Control Synthesis.
-Conformal prediction is an emerging technique in machine learning that provides a way to quantify the uncertainty of model predictions [18].
+Conformal prediction is an emerging technique in machine learning that provides a principle way to quantify the uncertainty in model predictions [18].
 It works by generating prediction sets or confidence regions that, with a specified level of certainty, are guaranteed to contain the true outcome.
@@
 However, a key distinction is that these approaches assume the behavior of uncontrollable agents is independent of the controlled system, an assumption that may not hold in many real-world scenarios.
-In contrast, our framework considers a coupled setting, addressing the issue of distribution shift in conformal prediction, where the behavior of uncontrollable agents is influenced by the state of the controlled system.
+In contrast, our framework considers a coupled setting, where the behavior of uncontrollable agents is influenced by the state of the controlled system.
 III Problem Formulation
-Notations: We denote by $\mathbb{R},\mathbb{N}$, and $\mathbb{R}^{n}$ the set of real numbers, natural numbers, and real vectors, respectively.
+Notations: We denote $\mathbb{R},\mathbb{N}$, and $\mathbb{R}^{n}$ as the set of real numbers, natural numbers, and real vectors, respectively.
 Let $\beta:\mathbb{R}\rightarrow\mathbb{R}$ denote an extended class $\mathcal{K}_{\infty}$ function, i.e., a strictly increasing function with $\beta(0)=0$.
-For a vector $v\in\mathbb{R}^{n}$, let $\|v\|$ denote its Euclidean norm.
-For a bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, let $\partial\mathcal{S}$ denote its boundary.
+For a vector $v\in\mathbb{R}^{n}$, let $|v|$, $\|v\|$, $\|v\|_{\infty}$ denote its $\ell_{1}$-norm, Euclidean norm and $\ell_{\infty}$-norm.
+For a finite set $D$, let $|D|$ denote its cardinality.
 III-A System Models
@@
 $\dot{x}(t)=f(x(t))+g(x(t))u(t)=:F(x(t),u(t)),$ (1)
-where $x(t)\in\mathbb{R}^{n}$ represents the state of the system at time $t$, and $u(t)\in\mathcal{U}$ denotes the control input, with $\mathcal{U}\subseteq\mathbb{R}^{m}$ being the set of admissible control inputs.
-The functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ represent the internal system dynamics and the input influence, respectively, and both are assumed to be locally Lipschitz continuous.
-For convenience, we denote $\mathbf{X}(t)=[x(t),\dot{x}(t)]\in\mathbb{R}^{n_{X}}$.
-Uncontrollable Agent: We consider the scenario where the controlled system in Eq.
-(1) operates among some uncontrollable agents.
-For the sake of simplicity, we only consider a single uncontrollable agent, and our approach can be extended to $N$ uncontrollable agents by considering their states as a vector.
-We denote the state of the agent at time instant $t$ by $Y(t)$, where $Y(t)\in\mathbb{R}^{n_{Y}}$.
-Note that $n_{Y}$ does not necessarily equal $n$, since the dynamics of $X$ and $Y$ may differ.
-More specifically, we assume that the uncontrollable agent has its own dynamic and decision law (intention), which can be either deterministic or stochastic.
-However, we do not assume to have any a priori knowledge of the agent dynamics or intentions.
-Note that, in general, the behavior of the uncontrollable agent depends on the state of the controlled agent.
-For example, in autonomous driving scenarios with moving pedestrians (uncontrollable agents), the behavior of the pedestrians is significantly influenced by the position and velocity of the car.
-Specifically, the closer the car is, the faster pedestrians tend to move away, or the faster the car approaches, the more cautious pedestrians become in their walking strategies.
-The objective is to synthesize a control input $u(t)$ for each time instant such that the controlled system can ensure safety with respect to the uncontrollable agent, with a certain level of confidence.
-Formally, safety here can be framed as a safe set $\mathcal{S}\subseteq\mathbb{R}^{n}\times\mathbb{R}^{n_{Y}}$.
-Then the problem formulation is as follows.
+where $x(t)\in\mathbb{R}^{n}$ represents the state of the system at time $t$, and $u(t)\in\mathcal{U}$ denotes the control input with $\mathcal{U}\subseteq\mathbb{R}^{c}$ being the set of admissible control inputs.
+The functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times c}$ represent the internal system dynamics and the input influence, respectively, and both are assumed to be locally Lipschitz continuous.
+Uncontrollable Agent: We consider a scenario where the controlled system is operating in an environment with a single uncontrollable agent for the sake of simplicity.
+And it can be easily extended to $N$ uncontrollable agents in principle by concatenating their states and considering them as a whole.
+Specifically, we assume that the behavior of the uncontrollable agent is coupled with the controllable agent, i.e., the evolution of the dynamics of uncontrollable agents is conditioned on the state of the controllable agent $x(t)$.
+For instance, in autonomous driving scenarios involving moving pedestrians, their behavior is strongly influenced by the car’s position and velocity.
+Specifically, pedestrians tend to move away more quickly as the car gets closer, and they adopt more cautious walking strategies when the car approaches at higher speeds.
+The uncontrollable agent dynamics can be modeled as a stochastic process
+$\dot{y}(t)\sim\mathbb{F}_{y}(x(t),y(t)),$ (2)
+where $y(t)\in\mathbb{R}^{m}$ is the state of the uncontrollable agent, and $\mathbb{F}_{y}$ is a distribution parameterized by the states of the controlled and uncontrollable systems, which is not known a priori.
+We frame safety as the 0-upper level set of a Lipschitz continuous function $h_{xy}(x,y)$, i.e., $\mathcal{S}=\{(x,y)\mid h_{xy}(x,y)\geq 0\}\subseteq\mathbb{R}^{n}\times\mathbb{R}^{m}$.
+And our objective is to synthesize a safety filter for reference control at each time instant such that the controlled system can ensure safety with respect to the uncontrollable agent, with a certain level of confidence.
 Problem 1
-Consider a controlled system described by Eq.
-(1) and an uncontrollable agent whose behavior depends on the behavior of the controlled system.
-Suppose that the controller can access the states of both the controlled system and uncontrollable agent precisely without measurement error, but the physical dynamics and intentions of the uncontrollable agent is unknown a priori.
-Design a controller that issues control input $u(t)$ for each time instant such that the closed-loop behavior is safe w.r.t.
-safety set $\mathcal{S}$, with a certain level of confidence.
-Since the behavior of the uncontrollable agent is complete unknown, hereafter, we will use sample data to provide confidence level guarantee for safety of form: “the probability that the system becomes unsafe within a certain time horizon is smaller than a threshold bound $\alpha$”.
+Given $h_{xy}$ defined with a safety set $\mathcal{S}$, a reference control $u_{r}$ at time $t$, we wish to obtain a safe control input $u^{*}$, that minimally interferes with the reference control while assuring safety.
+Formally, we aim to solve the following problem
+$u^{*}$ $=\arg\min_{u\in\mathcal{U}}||u-u_{r}||^{2}$ (3)
+s.t. $\text{Prob}[h_{xy}(x(t^{\prime}),y(t^{\prime}))\geq 0]\geq 1-\alpha,\;\forall t^{\prime}>t,$
+[…]
+[…]
+Since the behavior of the uncontrollable agent is complete unknown, we will train a neural network to model its behavior and provide confidence level guarantee for safety of form: “the probability that the system becomes unsafe with the filtered input is lower than a threshold $\alpha$”.
 IV Safe Prediction-Based Robust Controller for Coupled Agents (SPARC)
@@
 Before we formally present the implementation details of the proposed controller, we provide a high level overview of the entire SPARC framework.
-  • At each discrete time step, referred to as a prediction update point, we use a neural network to predict the intention of the uncontrollable agent over the next control period $\Delta t$.
-  • We apply the technique of conformal predictions to enlarge the predicted results into a region that contains the agents’ reachable states with high confidence.
+  • At each discrete time step, referred to as a prediction update point, we use a neural network to forecast the behavior of the uncontrollable agent over the next control period $\Delta t$.
+  • We apply conformal prediction techniques to expand the predicted results into a confidence region that contains the agent’s reachable states with high probability.
   • Then within each control period $\Delta t$, we apply the technique of control barrier functions to generate control inputs continuously to avoid reaching the predicted region of the uncontrollable agent.
@@
 We denote by $[X_{0},X_{1},\ldots,X_{k}]$ and $[Y_{0},Y_{1},\ldots,Y_{k}]$ the sampled trajectories.
-To capture the coupling between the controlled system and the uncontrollable agents, at each sampling instant, the movement of the uncontrollable agent, i.e., $\Delta Y_{i+1}:=Y_{i+1}-Y_{i}$, follows a distribution parameterized by the state of the controlled system, i.e.,
-$\Delta Y_{i+1}\sim\mathcal{D}(X_{i}),.$ (2)
-Note that, distribution $\mathcal{D}(X_{i})$ is memoryless and time-invariant meaning that the behavior of the uncontrollable agent only depends on the current state of the system rather than the entire history.
+For convenience, let $X$ denote the state space of the controllable agent which includes both position and velocity, while $Y$ represents the pedestrian’s position.
+To capture the coupling between the controlled system and the uncontrollable agents, at each sampling instant, the movement of the uncontrollable agent, i.e., $\Delta Y_{i}:=Y_{i+1}-Y_{i}$, follows a distribution parameterized by the state of the controlled system in (2), i.e.,
+$\Delta Y_{i}\sim\mathcal{D}(X_{i},Y_{i}).$ (4)
+Note that distribution $\mathcal{D}(X_{i},Y_{i})$ is memoryless and time-invariant meaning that the behavior of the uncontrollable agent only depends on the current state of the system rather than the entire history.
 This assumption is valid in many applications such as the behavior of moving pedestrians.
@@
 We assume that we have access to $\bar{K}$ independent trajectories of observations $X^{(j)}=(X_{1}^{(j)},X_{2}^{(j)},\ldots)$, $Y^{(j)}=(Y_{1}^{(j)},Y_{2}^{(j)},\ldots)$.
-Let $\Delta Y_{i+1}=Y_{i+1}-Y_{i}$, $\Delta Y^{(j)}=(\Delta Y_{1}^{(j)},\Delta Y_{2}^{(j)},\ldots),\forall j=1,2,\ldots,\bar{K}$ sampled from the distribution of $X^{(j)}$.
-Let $Z^{(j)}=(Z_{1}^{(j)},Z_{2}^{(j)},\ldots)$ where $Z_{i}^{(j)}=(X_{i}^{(j)},\Delta Y_{i}^{(j)})$.
-We partition these trajectories into calibration and training sets, denoted as $D_{\text{cal}}:=\{Z^{(1)},\dots,Z^{(Card(D_{cal}))}\}$ and $D_{\text{train}}:=\{Z^{(Card(D_{cal})+1)},\dots,Z^{(\bar{K})}\}$, respectively.
+Let $Z^{(j)}_{(0,t]}:=(Z_{1}^{(j)},\dots,Z_{t}^{(j)})$ where $Z_{i}^{(j)}=(X_{i}^{(j)},Y_{i}^{(j)})$.
+Let $\Delta Y^{(j)}=(\Delta Y_{1}^{(j)},\Delta Y_{2}^{(j)},\ldots),\forall j=1,2,\ldots,\bar{K}$ sampled from the random distribution of $(X^{(j)},Y^{(j)})$, we denote $W^{(j)}=[(Z^{(j)}_{(0,1]},\Delta Y^{(j)}_{1}),(Z^{(j)}_{(0,2]},\Delta Y^{(j)}_{2}),\dots]$
+We partition these trajectories into calibration and training sets, denoted as $D_{\text{cal}}:=\{W^{(1)},\dots,W^{(|D_{\text{cal}}|)}\}$ and $D_{\text{train}}:=\{W^{(|D_{\text{cal}}|+1)},\dots,W^{(\bar{K})}\}$, respectively.
 Trajectory Predictor.
 Our objective is to use the training set $D_{\text{train}}$ to design a predictor to forecast the future state of the uncontrollable agent.
-Note that, since $\Delta Y_{i+1}+Y_{i}=Y_{i+1}$, it suffices to predict $\Delta Y_{i+1}$ as we assume that all state information can be observed perfectly.
-In principle, although the $\Delta Y_{i+1}$ only depends on the current states, we still use the entire past samples to perform the prediction to improve the prediction accuracy.
+Note that, since $\Delta Y_{i}+Y_{i}=Y_{i+1}$, it suffices to predict $\Delta Y_{i}$ as we assume that all state information can be observed perfectly.
+In principle, although the $\Delta Y_{i}$ only depends on the current states, we still use the entire past samples to perform the prediction to improve the prediction accuracy considering the random disturbance influence on the uncontrollable agent dynamic.
 Therefore, the predictor is of form
-$\Delta\hat{Y}_{t+1}=\Omega(Z_{(0,t]}|\theta).$ (3)
-Note that $\Omega$ could be implemented by any trajectory predictor such as recurrent neural networks (RNNs), long short term memory (LSTM) networks or transformers , and $\theta$ is its parameters.
-IV-C Conformal Prediction for Reachable Region
-The predictions $\Delta\hat{Y}_{t}$, which may be generated by models with learning capabilities, are not guaranteed to be accurate.
-To address this, we employ conformal prediction to construct prediction regions for $\Delta\hat{Y}_{t}$ that hold with high probability.
-Lemma 1
-([25]).
-Let $V,V^{(1)},\ldots,V^{(k)}$ be $k+1$ independent and identically distributed real-valued random variables.
-Without loss of generality, let $V^{(1)},\ldots,V^{(k)}$ be sorted in non-decreasing order and define $V^{(k+1)}:=\infty$.
-For $\alpha\in(0,1)$, it holds that $\operatorname{Prob}(V\leq\bar{V})\geq 1-\alpha$ where
-$\bar{V}:=V^{(r)}\text{ with }r:=\lceil(k+1)(1-\alpha)\rceil$
-[…]
-In the context of our approach, we aim to use conformal prediction to quantify the uncertainty in the predicted state of the uncontrollable agent.
+$\Delta\hat{Y}_{t}=\Omega(Z_{(0,t]}|\theta).$ (5)
+Note that $\Omega$ could be implemented by any trajectory predictor such as neural networks (NNs) with parameters $\theta$.
+IV-C Uncertainty Qualification with Conformal Prediction
+The prediction $\Delta\hat{Y}_{t}$ generated by the trained model, is not guaranteed to be accurate.
+To address this, we employ conformal prediction to construct a prediction region for $\Delta\hat{Y}_{t}$ that holds with high probability.
 Definition 1
-For a trajectory $Z^{(j)}=(Z_{1}^{(j)},Z_{2}^{(j)},\ldots)\in D_{cal}$ and $Z^{(j)}_{i}=(X^{(j)}_{i},\Delta Y^{(j)}_{i})$, the prediction error is then defined as
-$R^{(j)}:=\sup_{Z^{(j)}_{i}\in Z^{(j)}}||\Delta Y_{i}^{(j)}-\Delta\hat{Y}_{i}^{(j)}||.$ (4)
-$\Delta Y_{i}^{(j)}$ is regarded as the ground truth of $\Delta\hat{Y}_{i}^{(j)}$.
+For a joint trajectory $Z^{(j)}_{(0,i]}=(Z_{1}^{(j)},Z_{2}^{(j)},\ldots,Z_{i}^{j})$ and $W^{(j)}_{i}=(Z^{(j)}_{(0,i]},\Delta Y^{(j)}_{i})\in D_{\text{cal}}$, the nonconformity score $R^{(j)}$ is defined as the prediction error
+$R^{(j)}:=\sup_{W^{(j)}_{i}\in W^{(j)}}||\Delta Y_{i}^{(j)}-\Delta\hat{Y}_{i}^{(j)}||,$ (6)
+where $\Delta Y_{i}^{(j)}$ is the ground truth of a predicted $\Delta\hat{Y}_{i}^{(j)}$.
 The following result states how to use the calibration dataset to generate a confidence regions that account for potential variations in the agent’s behavior.
 Proposition 1
-Given a calibration dataset $D_{cal}\subseteq\mathbb{R}^{(n_{X},n_{Y}})$, and a failure probability tolerance $\alpha\in[0,1]$, by Lemma 1, $\forall X\in\mathbb{R}^{n_{X}}$, $t\in\mathbb{R}_{\geq 0}$, $\operatorname{Prob}\left(R^{(0)}\leq\bar{R}(\alpha)\right)\geq 1-\alpha$ holds where
-$\bar{R}(\alpha)=\text{ Quantile }_{1-\alpha}\left(R^{(1)},\ldots,R^{(Card(D_{cal}))},\infty\right)$ (5)
-Note that here $\operatorname{Card}(D_{\text{cal}})$ is the dataset size as in Assumption 1.
-Similarly, we obtain $\bar{R}=R^{(r)}$, where $r=\lceil\operatorname{Card}(D_{\text{cal}})\cdot(1-\alpha)\rceil$.
-As before, for meaningful prediction regions, we require $\operatorname{Card}(D_{\text{cal}})\geq\lceil(\operatorname{Card}(D_{\text{cal}})+1)(1-\alpha)\rceil$, otherwise $\bar{R}(\alpha)=\infty$.
-Proof:
-In Lemma 1, given a set of independent and identically distributed random variables $V^{(1)},V^{(2)},\ldots,V^{(k)}$, we can derive a bound in a sense of probability, by sorting these values and determining the corresponding quantile.
-In Proposition 1, we treat the nonconformity scores $R^{(1)},R^{(2)},\ldots,R^{(\text{Card}(D_{\text{cal}}))}$ as analogous to the random variables $V^{(1)},V^{(2)},\ldots,V^{(k)}$ in Lemma 1.
-Therefore, $R$ plays the role of $V$ in the lemma, allowing us to directly apply its result.
-∎
-The above result says that the state estimation error $R$ can essentially be bounded, with probability $1-\alpha$, by the worst case of conformal prediction region $\bar{R}(\alpha)$.
-Since the score $R^{(0)}$ represents the model’s estimation error sampled from any trajectory that might occur during the actual operation of the system, we can hence conclude that
-$\operatorname{Prob}[||\Delta Y_{t+1}-\Delta\hat{Y}_{t+1}||\leq\bar{R}(\alpha)]\geq 1-\alpha$ (6)
-From known measurement $X_{t}$ and $Y_{t}$, and prediction $\Delta\hat{Y}_{t+1}$, denote safe region in a probabilistic sense for $\Delta\hat{Y}_{t}$ and $\bar{R}(\alpha)$ as $\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)=[(Y_{t}+\Delta\hat{Y}_{t+1})-\bar{R}(\alpha),(Y_{t}+\Delta\hat{Y}_{t+1})+\bar{R}(\alpha)]\subseteq\mathbb{R}^{n_{Y}}$ satisfies that
-$\operatorname{Prob}[Y_{t+1}\in\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)]\geq 1-\alpha$ (7)
-Note that $\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)$ is the estimated safe region of the uncontrollable event within the next sample period.
+Given a calibration dataset $D_{\text{cal}}$, and a failure probability tolerance $\alpha\in[0,1]$, by [25], $\forall X\in\mathbb{R}^{n_{X}}$, $t\in\mathbb{R}_{\geq 0}$, $\operatorname{Prob}\left(R^{(0)}\leq\bar{R}(\alpha)\right)\geq 1-\alpha$ holds where
+$\bar{R}(\alpha)=\text{ Quantile }_{1-\alpha}\left(R^{(1)},\ldots,R^{(|D_{\text{cal}}|)},\infty\right)$ (7)
+Note that here $\bar{R}(\alpha)\in\mathbb{R}^{m}$ is a vector for each dimension and $|D_{\text{cal}}|$ is the calibration dataset size.
+Similarly, we obtain $\bar{R}=R^{(r)}$, where $r=\lceil|D_{\text{cal}}|\cdot(1-\alpha)\rceil$.
+As before, for meaningful prediction regions, we require $|D_{\text{cal}}|\geq\lceil(|D_{\text{cal}}|+1)(1-\alpha)\rceil$, otherwise $\bar{R}(\alpha)=\infty$.
+Conformal prediction implies the nonconformity score $R$ can essentially be bounded, with probability $1-\alpha$, within the conformal region $\bar{R}(\alpha)$.
+Given that the test and calibration data are exchangeable, we can hence conclude that for each coordinate direction $i=1,...,m,$
+$\operatorname{Prob}\left[\left(\frac{|y_{i}-(Y_{t}+\Delta\hat{Y}_{t})_{i}|}{\bar{R}_{i}(\alpha)}\right)\leq 1\right]\geq 1-\alpha$ (8)
+From known measurement $X_{t}$ and $Y_{t}$, and prediction $\Delta\hat{Y}_{t}$, we denote the conformal region for $Y_{t+1}$ in a probabilistic sense as
+$\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)=\left\{y\in\mathbb{R}^{m}\mid\max_{i}\left(\frac{|y_{i}-(Y_{t}+\Delta\hat{Y}_{t})_{i}|}{\bar{R}_{i}(\alpha)}\right)\leq 1\right\}$ (9)
+This set represents a weighted $\ell_{\infty}$-norm ball centered at $(Y_{t}+\Delta\hat{Y}_{t})$, where $\bar{R}(\alpha)$ determines the scaling in each coordinate direction. It satisfies that
+$\operatorname{Prob}[Y_{t+1}\in\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)]\geq 1-\alpha.$ (10)
+Note that $\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)$ is the estimated conformal prediction region of the uncontrollable agent in the next sampling step.
 IV-D Safe Control using Control Barrier Functions
-After bounding the state estimation error in Proposition 1, we calculated the conformal prediction region bound $\bar{R}$ for Eq.
-(6) to design a uncertainty-aware controller via the technique of control barrier functions.
+After we quantified the uncertainty with the calculated prediction region, we design an uncertainty-aware safety filter with CBF.
 Definition 2
@@
 where $L_{f}h(x)$ and $L_{g}h(x)$ represent the Lie derivatives.
+At each time step, the state of the uncontrollable agent is treated as a fixed parameter, reducing $h_{xy}(x,y)$ to $h(x)$.
 Therefore, Problem 1 can be formulated as a safe control problem within each prediction update period by considering the estimated region of the uncontrollable agent as an obstacle simply by define
-$\mathcal{C}=\{x\in\mathbb{R}^{n}\mid\forall Y\in\mathbb{R}^{n_{Y}}:(x,Y)\in\mathcal{S}\}$ (9)
-as the safe set for the controlled system such that the overall system remain in safe set $\mathcal{S}$. Therefore, it suffices to find a CBF $\hat{h}(x)$ such that
-$\hat{C}=\{x\in\mathbb{R}^{n}\mid\hat{h}(x)\geq 0\}\subseteq\mathcal{C}.$ (10)
-In the literature, there are many existing approach for finding such CBF $\hat{h}(x)$, e.g., by sum-of-square programming [26] or learning-based approach [27, 28].
-Now, given the synthesized CBF $\hat{h}(x)$, the set of admissible control inputs is defined by
-$K_{CBF}:=\{u\in\mathcal{U}\mid L_{f}\hat{h}(x)+L_{g}\hat{h}(x)u\geq-\beta(\hat{h}(x))\}$ (11)
-The following result shows that, by following the control strategy in Eq. (11), one can ensure safety within each period $\Delta$ with confidence $1-\alpha$.
+$\mathcal{C}=\{x\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{m}:(x,y)\in\mathcal{S}\}$ (12)
+as the safe set for the controlled system such that the overall system remain in safe set $\mathcal{S}$. Therefore, it suffices to find a CBF $h(x)$ such that
+$C=\{x\in\mathbb{R}^{n}\mid{h}(x)\geq 0\}\subseteq\mathcal{C}.$ (13)
+In the literature, there are many existing approach for finding such CBF $h(x)$, e.g., by sum-of-square programming [26] or learning-based approach [27, 28].
+Now, given the synthesized CBF $h(x)$, the set of admissible control inputs is defined by
+$K_{CBF}:=\{u\in\mathcal{U}\mid L_{f}h(x)+L_{g}h(x)u\geq-\beta(h(x))\}$ (14)
+The safe input can be any control law $u(t)$ from $K_{CBF}$ in Eq. (14), e.g., one computed by solving the quadratic optimization problem (3).
 Proposition 2
-Consider the control law $u(t)$ from $K_{CBF}$ in Eq.
-(11),the system composed by controllable and uncontrollable variables Eq.
-(1), Eq.
-(2) satisfies Eq.
-(8) where $\mathcal{L}_{L_{f}h},\mathcal{L}_{\beta\circ h}$, and $\mathcal{L}_{L_{g}h}$ are the Lipschitz constants of the functions $L_{f}h,\beta\circ h$ and $L_{g}h$, respectively, we have
+Consider the control law $u(t)$ from $K_{CBF}$ in Eq. (14),the system composed by controllable and uncontrollable variables Eq. (1), Eq. (4) satisfies Eq. (11), respectively, we have
 $\operatorname{Prob}[(X_{t+1},Y_{t+1})\in\mathcal{S}]\geq 1-\alpha$ (15)
 Proof:
-When $Y_{t+1}\in\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)$ (with probability higher than $1-\alpha$), the“obstacle”(i.e., the safe region which includes the actual position of the uncontrollable agent inside) is accurately estimated.
-Consequently, the control law $u(t)$ derived from $K_{CBF}$ for $\hat{\mathcal{C}}$ is effective in ensuring the safety of the entire system, meaning that $(X_{t+1},Y_{t+1})$ remains within the safe set $\mathcal{S}$, as defined in Problem 1.
+With guarantee that $\operatorname{Prob}[Y_{t+1}\in\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)]\geq 1-\alpha$, the control law $u(t)$ derived from $K_{CBF}$ for $\hat{\mathcal{C}}$ is effective in ensuring the safety of the entire system, meaning that $(X_{t+1},Y_{t+1})$ remains within the safe set $\mathcal{S}$, as defined in Problem 1.
 ∎
@@
 In this section, we illustrate and evaluate the proposed approach through simulations of a case study involving autonomous driving with uncontrollable pedestrians.
-This case study demonstrates the effectiveness of our framework in managing the interactions between the autonomous vehicle and pedestrians, ensuring safety despite the uncertainty in pedestrian behavior.
-Figure 3: The pedestrian’s walking logic is modeled as an unknown stochastic distribution affected by the state of the vehicle, i.e.,$\Delta Y\sim\mathcal{D}(X)$.
+This case study demonstrates the effectiveness of our framework in modelling the interactions between the autonomous vehicle and pedestrians, providing probabilistic safety guarantees under the uncertainty of pedestrian behavior.
+Figure 3: The pedestrian’s speed follows an unknown stochastic distribution affected by the state of the vehicle, i.e.,$\Delta Y\sim\mathcal{D}(X,Y)$.
 The pedestrian starts at a random position on one side of the crosswalk while the vehicle begins at the midpoint of the starting line.
 Red circles indicate collision volume.
-Figure 4: Nonconformity scores histogram for pedestrian position displacement $(\Delta x,\Delta y)$ on $\mathcal{D}_{cal}$
-V-A Scenario Description and System Dynamic
+Figure 4: Nonconformity scores histogram for pedestrian position displacement $(\Delta{Y_{par}},\Delta{Y_{perp}})$ on $D_{cal}$
+V-A Scenario Description and System Dynamics
 Scenario Setup: As illustrated in Fig.
@@
 The control period is set to be $\Delta t$ = 0.1s.
-Vehicle Dynamics: The vehicle’s starting speed is sampled from a uniform distribution within the range $[0,15]$ m/s, which mimics typical vehicle speeds observed in urban environments (up to the speed limit of 15 m/s).
+Vehicle Dynamics: For the sake of simplicity, we assume the vehicle moves straight through the crosswalk, setting the dimension $n=1$ with dynamic model $\dot{x}=u$.
+This choice aligns with real-world constraints, as vehicles rarely make left or right turns while crossing a zebra line.
+Nevertheless, our approach naturally extends to two or higher dimensions.
+The vehicle’s speed range is $[0,15]$m/s, reflecting typical vehicle speeds observed in urban environments (up to the speed limit of 15 m·s-1/54 km·h-1/33 mph).
 The vehicle’s position is updated based on its velocity.
-Pedestrian Behavior: The pedestrian’s movement is modeled as a unknown distribution of the vehicle’s state-space vector and Gaussian noise i.e.
-$\Delta{Y}=\phi(X)+\mathcal{N}({0,\sigma})\sim\mathcal{D}(X)$ (13)
-[…]
+Pedestrian Behavior: The pedestrian’s movement is modeled as a stochastic process governed by a function of the vehicle’s state-space vector and Gaussian noise.
+Mathematically, this can be expressed as:
+$\Delta{Y}\sim\mathcal{N}_{2}(\phi(X,Y),\Sigma),$ (16)
+where $\Delta Y=(\Delta{Y_{par}},\Delta{Y_{perp}})\in\mathbb{R}^{2}$ follows a Gaussian distribution whose mean value is a function of both the state of the pedestrian and the controlled vehicle.
+Here the notation $\Delta{Y_{par}}$ represents the velocity component parallel to the road direction, while $\Delta{Y_{perp}}$ denotes perpendicular component.
+$\Sigma$ is the standard deviation $\Sigma=\operatorname{diag}(\sigma^{2}_{x},\sigma^{2}_{y})=\operatorname{diag}((\text{0.5m/s})^{2},(\text{2m/s})^{2})$.
+To illustrate, as the vehicle approaches at a higher speed, the pedestrian slows down $\Delta{Y_{perp}}$, exhibiting more cautious crossing behavior.
+Simultaneously, $\Delta{Y_{par}}$ increases in the direction of moving away from the vehicle as an evasive response.
+Additionally, $\Delta{Y_{par}}$ is influenced by $Y$.
+When approaching the left/right edge of the crosswalk area, the pedestrian tends to adjust the position toward the center of the crosswalk.
+And as pedestrian movement is inherently unpredictable, this modelling introduces additional stochasticity to account for hesitation and spontaneous decision-making.
+Under this simulation setup, pedestrians exhibit a perturbed yet generally arced crossing trajectory around the vehicle.
+This behavior aligns with the natural tendency of pedestrians to adjust their paths dynamically while crossing.
+V-B Model Implementation
 Interaction and Collision Avoidance: The pedestrian’s trajectory is evaluated in real-time as the vehicle approaches.
-We track the pedestrian’s crossing behavior when the pedestrian is on the crosswalk, where the vehicle’s speed directly impacts the pedestrian’s decision to speed up, slow down, or stop.
+We track the pedestrian’s crossing behavior which is influenced by the vehicle.
 The simulation continuously monitors the distance between the vehicle and the pedestrian, and the safe controller (SPARC) is applied to ensure safe control synthesis in avoiding collisions.
 Trajectory Predictor: We trained a feedforward neural network to predict the future trajectory of the pedestrian.
+The dataset is generated by simulating a vehicle moving at a constant speed while tracking both the vehicle and pedestrian trajectories.
+The vehicle’s speed is uniformly sampled from the range $[0,15]$m/s.
+The collected data is structured following the format described in the IV.
+B section.
 The input is the history state measurement of the vehicle’s and pedestrian’s state.
-The training dataset $D_{\text{train }}$ contains $1\times 10^{6}$ data points, and the calibration dataset $D_{\text{cal }}$ for conformal prediction contains $1\times 10^{5}$ data points.
-Barrier Functions: For vehicle-pedestrian collision avoidance, we may define the CBF as follows:
-$h(x)=\min_{Y\in\partial\hat{\mathcal{O}}_{Y}}||x-Y||-d_{\text{safe}}$ (14)
+The training dataset $D_{\text{train }}$ contains $1\times 10^{6}$ data, and the calibration dataset $D_{\text{cal}}$ for conformal prediction contains $1\times 10^{5}$ data.
+The model consists of three fully connected layers with ReLU activation.
+The input layer maps features to 64 hidden units, followed by a second layer of 64 units, and a final output layer predicting pedestrian speed $(\Delta{Y_{par}},\Delta{Y_{perp}})\in\mathbb{R}^{2}$.
+The model use the Adam optimizer (learning rate 0.001) and MSE loss.
+A StepLR scheduler halves the learning rate every 10 epochs over 50 epochs.
+Control Barrier Functions: For vehicle-pedestrian collision avoidance, we choose the following CBF:
+$h(x)=\min_{y\in\hat{\mathcal{O}}_{y}}||x-y||-d_{\text{safe}}$ (17)
 where:
-  • $x\in\mathbb{R}^{n}$ is the position of the vehicle
-  • $\partial\hat{\mathcal{O}}_{Y}\subseteq\mathbb{R}^{n_{Y}}$ is the predicted safe region boundary of the uncontrollable pedestrian
-  • $d_{\text{safe}}$ is the minimum safe distance to be maintained
+  • $x\in\mathbb{R}^{n}$ is the position of the vehicle,
+  • $\hat{\mathcal{O}}_{y}\subseteq\mathbb{R}^{m}$ is the predicted conformal region of the uncontrollable pedestrian, and we set $\hat{\mathcal{O}}_{y}(\alpha)=[\hat{y}-\bar{R}(\alpha),\hat{y}+\bar{R}(\alpha)]\subseteq\mathbb{R}^{2}$ as a weighted $\ell_{\infty}$-norm ball.
+  • $d_{\text{safe}}$ is the minimum safe distance to be maintained,
 […]
-This CBF in a sense of probability ensures a safe distance is maintained between the vehicle and the pedestrian.
-The system is in a safe state, avoiding collision, when $h(x)\geq 0$.
-The controller design should satisfy the inequality in Eq.
-(8) to guarantee system safety.
-Set $d_{\text{safe}}$ = 1.0 m, the collision volume is shown as the red circle in Fig.
-3.
+With $d_{\text{safe}}$ = 1.0 m, the collision volume is shown as the red circle in Fig. 3.
 To demonstrate the effectiveness of our method, we compare the following two cases in simulations:
   1. 1.
-    Uncontrollable Prediction Robust CBF: Let $\alpha$ be the failure tolerance, randomly pick $\alpha=0.85,0.75,0.5$, the results are shown in Table I.
+    SPARC: Let $\alpha$ be the failure tolerance, randomly pick $\alpha=0.85,0.75,0.5$, the results are shown in Table I.
   2. 2.
-    Vanilla CBF: We set $\alpha=0\implies\bar{R}(\alpha)=[0,0,\ldots,0]$, which essentially reduces to the vanilla CBF [8].
-The barrier function is directly from neural network output rather than safe region, so this CBF cannot provide any safety guarantee.
-V-B Result Analysis
-Figure 5: The result of the conformal prediction in the case study
-Trajectory Predictors and Conformal Prediction: Take failure probability tolerance $\alpha=0.85$ as an example, Fig.
-4 illustrates a calculation of the nonconformity scores used in our approach.
-The scores $R_{x}^{0.85}=0.75,R_{y}^{0.85}=2.86$ represent the deviation between the predicted pedestrian position change $\Delta\hat{Y}=(\Delta x,\Delta y)$ and the ground truth, providing a measure of prediction uncertainty , which is used to generate safe control decisions.
-Fig.
-5 shows the network successfully learns the underlying random distribution governing pedestrian behavior in relation to the vehicle’s state.
-In our simulation framework, for the sake of simplicity, the upper limit of all nonconformity scores is employed as a proxy for the overall safety margin, and it is also feasible to utilize a grid-based distribution of nonconformity scores to devise a more aggressive control strategy.
-Safe Control Barrier Function Controller: The vehicle driven by our SPARC is expected to avoid collisions with pedestrians.
-We set the failure probability tolerance $\alpha=0.15,0.25,0.5$, and desire $\text{Prob}[(X,Y)\in\mathcal{S}]\geq 1-\alpha$, which means $\text{Prob}[\text{vehicle and pedestrian collide}]\leq\alpha$.
-We performed $1\times 10^{5}$ experiments, each time step is set to 100, the collision probability are collected in Table I.
-In 11.20% of the trials(1120 experiments), the car encounter a pedestrian, and we care about whether our controller can successfully prevent collision in these experiment, so we divide absolute collision frequency by 11.20% to get relative collision frequency.
-The collision rate of SPARC are $14.46\%(<15\%),17.50\%(<25\%),37.59\%(<50\%)$, which are all under their respective failure probability tolerance and significantly lower than vanilla CBF case $(96.43\%)$, so our Proposition 2.
-holds in practice.
+    Vanilla CBF: We set $\alpha=1$, which means that $\bar{R}(\alpha)=[0,0]$.
+This essentially reduces to the vanilla CBF [8].
+This method uses the direct output from the neural network output than the calibrated conformal region, providing no safety guarantee and failing to account for dynamic stochasticity.
+  3. 3.
+    Random: In this experiment, the vehicle’s initial velocity is uniformly sampled, and it follows the reference control $u_{r}$ during its motion.
+The proportion of trials resulting in a collision with the pedestrian is recorded.
 TABLE I: Collision Probability
@@
 Random 11.20% 100.00%
+Figure 5: The prediction output with conformal prediction region. The plot also includes $\phi(\text{car speed})$ with all other parameters fixed as a noise-free reference ground truth.
+V-C Result Analysis
+Trajectory Predictors and Conformal Prediction: Take failure probability tolerance $\alpha=0.85$ as an example, Fig.
+4 illustrates a calculation of the nonconformity scores used in our approach.
+The scores $R_{Y_{par}}^{0.85}=0.75,R_{Y_{perp}}^{0.85}=2.86$ represent the deviation between the predicted pedestrian position change $\Delta\hat{Y}$ and the ground truth, providing a measure of prediction uncertainty , which is used to generate safe control decisions.
+Fig.
+5 shows the how the network learns the underlying random distribution governing pedestrian behavior in relation to the vehicle’s state.
+Notably, our method does not rely on the precise accuracy, overall performance, or scene comprehension capabilities of the model, making it robust to model imperfections and uncertainties.
+In our simulation framework, for the sake of simplicity, the upper limit of all nonconformity scores is employed as a proxy for the overall safety margin, and it is also feasible to utilize a grid-based distribution of nonconformity scores to devise a more aggressive control strategy.
+Safe Control Barrier Function Controller: The vehicle controlled by our SPARC framework is expected to avoid collisions with pedestrians.
+We set the failure probability tolerance $\alpha=0.15,0.25,0.5$, and desire $\text{Prob}[(X,Y)\in\mathcal{S}]\geq 1-\alpha$, which means $\text{Prob}[\text{vehicle and pedestrian collide}]\leq\alpha$.
+We performed $1\times 10^{5}$ experiments, each time step is set to 100, the collision probability are collected in Table I.
+In 11.20% of the trials (1120 experiments), the car crash on the pedestrian, and we care about whether our controller can successfully prevent collision in these experiment, so we divide absolute collision frequency by 11.20% to get relative collision frequency.
+Due to the relatively high ratio of the standard deviation of the added stochastic component to the pedestrian’s average speed, the vanilla CBF, which relies solely on predicted pedestrian positions for avoidance, struggles to account for this uncertainty and consequently suffers from a high failure rate$(96.43\%)$.
+The collision rate of SPARC are $14.46\%(<15\%),17.50\%(<25\%),37.59\%(<50\%)$, which are all under their respective failure probability tolerance and significantly lower than vanilla CBF case $(96.43\%)$, so our Proposition 2.
+holds in practice.
 VI Conclusion
@@
 Compared to previous approaches, our method accounts for the coupling between the controlled system and the uncontrollable agent by sampling and evaluating over an augmented state space, making it more applicable to real-world scenarios such as autonomous driving and human-robot interaction.
-However, one limitation of our approach is the assumption that precise state information for both the controlled system and uncontrollable agent is fully accessible.
+There are several future directions for this work that we plan to explore.
+The main objective of our work is to present a generic framework applicable to a broad class of control synthesis problems involving coupled uncontrollable agents, rather than being tailored to specific applications.
+In this work, for illustrative purposes, we use simplified vehicle and pedestrian models to demonstrate our approach.
+Regarding the complex nature of vehicle-pedestrian interactions, various datasets and prediction models have been developed to capture realistic behaviors.
+Notable datasets such as ETH & UCY [29] and JAAD [30] provide benchmarks for pedestrian trajectory modeling, while methods like Social-LSTM [31] and Trajectron++ [32] offer deep learning-based approaches for trajectory prediction.
+The TrajNet++ benchmark [33] further standardizes evaluation in this domain.
+In the future, we plan to focus specifically on autonomous driving by leveraging these real-world datasets and customized prediction methods to enhance the accuracy and applicability of our framework.
+Additionally, one limitation of our current approach is the assumption that precise state information for both the controlled system and the uncontrollable agent is fully accessible.
 In many practical applications, such state information must be obtained through perception devices like LiDAR, which can introduce estimation errors.
-In the future, we plan to extend our framework to perception-based settings, addressing the additional challenge of state estimation uncertainty.
+To address this, we plan to extend our framework to perception-based settings, incorporating techniques to handle state estimation uncertainty and further improve the robustness of our approach in real-world scenarios.